Hi all, I am still struggling a bit with this vectors stuff, hopefully I willahve it sorted after learning how to solve the following:
I need to calculate 2~u + ~v - 5~w, if ~u = 3~i + 2~j - 7~k, ~v = 5~i - 2~j + 4~k and ~w = 3~i + 2~j - 4~k.
(Agian the ~ symbols refer to arrows in the positive direction above the values shown)
Thnaks, -nick.

Answer by Alan3354(21583)

I need to calculate 2~u + ~v - 5~w, if ~u = 3~i + 2~j - 7~k, ~v = 5~i - 2~j + 4~k and ~w = 3~i + 2~j - 4~k.
-----------
The ~ don't help, they make it harder to see what you mean.
-----------
2u + v - 5w
if u = 3i + 2j - 7k, v = 5i - 2j + 4k and w = 3i + 2j - 4k
-------
+2u = 6i + 4j - 14k
+1v = 5i - 2j + 4k
-5w =-15i-10j + 20k
-------------------- Add them up
Sum = -4i -8j + 10k